### v1

@@ 4.4 Comparison with other Baselines
Mip-NeRF [1] 8.68 16.54 23.58 0.571 0.741 0.879 0.353 0.198 0.092 0.323 0.148 0.056
DietNeRF [9] 11.85 20.63 23.83 0.633 0.778 0.823 0.314 0.201 0.173 0.243 0.101 0.068
@@
${}^{\ddagger}$ [45] 19.92 23.25 25.38 0.781 0.838 0.877 0.125 0.085 0.057 0.086 0.058 0.038
@@
HourglassNeRF 19.85 22.73 25.14 0.773 0.842 0.886 0.146 0.084 0.057 0.096 0.060 0.040
mip-NeRF [1] MixNeRF [34] FreeNeRF [45] FlipNeRF [33] HourglassNeRF Ground Truth
(a) Realistic Synthetic ${}^{\circ}$ 4-view
(b) Realistic Synthetic ${}^{\circ}$ 8-view
mip-NeRF [1] RegNeRF [25] ${}^{\ddagger}$ [45] FlipNeRF [33] HourglassNeRF Ground Truth
(c) DTU 3-view
(d) DTU 6-view
(e) DTU 9-view
Figure 7: Qualitative comparisons on Realistic Synthetic ${}^{\circ}$ and DTU. Our HourglassNeRF achieves notable rendering quality with better capturing fine details. normal-$\ddagger$ indicates the W&B prior.
DTU.
Tab. 4 shows the quantitative comparisons on DTU. Our HourglassNeRF outperforms the baseline FlipNeRF, and achieves competitive results overall. As demonstrated in Fig. 7, ours shows notable outcomes with sharply rendered textures. FreeNeRF is trained with the black and white prior assuming the estimated black and white color as the background and table, respectively, which is a highly strong assumption specific to the dataset, and achieves degenerate results without the prior. However, our HourglassNeRF utilizes the Lambertian assumption, which is a strong but effective condition for few-shot scenario, reducing the heuristic components for the training framework. Furthermore, as shown in Fig. 9, while FreeNeRF achieves high-quality rendering, it relies on a white and black prior, i.e. a highly heuristic approach given the dataset’s characteristics. In contrast, our HourglassNeRF shows competitive performance without relying on such priors.
Figure 8: Additional comparison with FreeNeRF on DTU 3-view. Although FreeNeRF achieves high-quality of rendering with only a few images, it depends on the white and black prior, which is highly heuristic based on the characteristics of dataset. Our HourglassNeRF achieves comparable performance to FreeNeRF without the heuristic prior for training. normal-$\ddagger$ denotes the W&B prior.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average $\downarrow$ ${}^{\ddagger}$ [45] 19.92 0.781 0.125 0.086 FreeNeRF [45] 19.23 0.769 0.149 0.103 HourglassNeRF 19.85 0.773 0.146 0.096
${}^{\ddagger}$ [45] FreeNeRF [45] HourglassNeRF
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Average $\downarrow$ Ref-NeRF [40] 17.10 0.821 0.190 0.142 FreeNeRF [45] 16.99 0.828 0.157 0.131 FlipNeRF [33] 18.14 0.847 0.141 0.109 HourglassNeRF 18.68 0.851 0.141 0.107
FreeNeRF [45] FlipNeRF [33] Ours GT
Figure 8: Additional comparison with FreeNeRF on DTU 3-view. Although FreeNeRF achieves high-quality of rendering with only a few images, it depends on the white and black prior, which is highly heuristic based on the characteristics of dataset. Our HourglassNeRF achieves comparable performance to FreeNeRF without the heuristic prior for training. normal-$\ddagger$ denotes the W&B prior.
Figure 9: Quantitative and qualitative results on Shiny Blender 4-view. While FlipNeRF demonstrates competitive performance by utilizing flipped specular reflection rays, which are physically aligned with the non-Lambertian surfaces, our approach consistently outperforms FlipNeRF and other comparative methods.
Shiny Blender.
We additionally compare our HourglassNeRF and other methods on Shiny Blender, which is largely non-Lambertian, as in Fig. 9. Although FlipNeRF achieves competitive performance thanks to its ray augmentation strategy utilizing specular reflection, which accords with the non-Lambertian surfaces, ours still outperforms other methods including FlipNeRF. We conjecture that under the few-shot setting where the major challenge is to learn 3D geometry effectively while appropriately avoiding the overfitting, our high-frequency regularization effect plays an important role despite contradicting physical assumptions.
5 Conclusion
In this work, we have approached the few-shot novel view synthesis from the perspective of a novel ray parameterization. Our HourglassNeRF casts an hourglass as an augmented ray, which adaptively regularizes the high-frequency components based on the target pixel photo-consistency. Through our proposed hourglass casting strategy, our HourglassNeRF effectively prevents overfitting to high-frequency components while achieving sharp fine details. Furthermore, compared to the baseline FlipNeRF, ours is a more physically grounded training framework by conceptualizing our hourglass as a featurized bundle of flipped diffuse reflection rays with our proposed luminance consistency regularization. As the hourglass covers the wider area of unseen views than a single ray, our HourglassNeRF is trained more efficiently without casting multiple augmented rays per original ray. Our HourglassNeRF achieves the SOTA or competitive performance compared to other methods under various few-shot scenarios. We expect that our HourglassNeRF is able to inspire a meaningful approach in the few-shot NeRF framework.
Limitations and future work.
Since our HourglassNeRF is based upon the Lambertian assumption, we regard all reflection rays as diffuse even if they are from a shiny surface. As a result, although our HourglassNeRF shows superior performance to other methods on the Shiny Blender dataset mainly consisting of non-Lambertian surfaces, the characteristics of the object’s surface are not adequately considered. Using both specular and diffuse reflection rays adaptively based on the estimated texture of the surface can be a meaningful direction for future work.

### v2

@@ C Additional Experiments
Mip-NeRF [1] - 8.68 16.54 23.58 0.571 0.741 0.879 0.353 0.198 0.092 0.323 0.148 0.056
3DGS [14] 14.18 - - 0.628 - - 0.301 - - 0.191 - -
PixelNeRF [51] Pre-training 16.82 19.11 20.40 0.695 0.745 0.768 0.270 0.232 0.220 0.147 0.115 0.100
PixelNeRF [51] 18.95 20.56 21.83 0.710 0.753 0.781 0.269 0.223 0.203 0.125 0.104 0.090
SRF [5] 15.32 17.54 18.35 0.671 0.730 0.752 0.304 0.250 0.232 0.171 0.132 0.120
SRF [5] 15.68 18.87 20.75 0.698 0.757 0.785 0.281 0.225 0.205 0.162 0.114 0.093
MVSNeRF [4] 18.63 20.70 22.40 0.769 0.823 0.853 0.197 0.156 0.135 0.113 0.088 0.068
MVSNeRF [4] 18.54 20.49 22.22 0.769 0.822 0.853 0.197 0.155 0.135 0.113 0.089 0.069
DietNeRF [9] Regularization 11.85 20.63 23.83 0.633 0.778 0.823 0.314 0.201 0.173 0.243 0.101 0.068
@@
SimpleNeRF [36] 16.25 20.60 22.75 0.751 0.828 0.856 0.249 0.190 0.176 0.143 0.088 0.071
DiffusioNeRF [47] 16.20 20.34 25.18 0.698 0.818 0.883 0.160 0.093 0.046 0.128 0.072 0.036
SparseNeRF [43] 19.55 - - 0.769 - - 0.201 - - 0.102 - -
FreeNeRF [50] 19.92 23.25 25.60 0.781 0.838 0.877 0.125 0.085 0.057 0.086 0.058 0.038
FreeNeRF [50] 19.23 22.77 25.59 0.769 0.835 0.877 0.149 0.088 0.057 0.103 0.063 0.039
@@
SparseGS [48] 18.89 - - 0.702 - - 0.229 - - 0.117 - -
ARC-NeRF 19.85 22.73 25.14 0.773 0.842 0.886 0.146 0.084 0.057 0.096 0.060 0.040
Table D: Additional quantitative comparison on DTU. Our ARC-NeRF shows competitive performance by outperforming other methods on most metrics, even without relying on any dataset-specific priors. $\dagger$ and $\ddagger$ indicate fine-tuning and W&B prior, respectively.
mip-NeRF [1] MixNeRF [35] FreeNeRF [50] FlipNeRF [34] ARC-NeRF Ground Truth
(a) Realistic Synthetic 360 8-view
mip-NeRF [1] RegNeRF [26] FreeNeRF [50] FlipNeRF [34] ARC-NeRF Ground Truth
(b) DTU 6-view
(c) DTU 9-view
Figure C: Additional qualitative comparisons. Our ARC-NeRF achieves high-quality renderings with fine details and clearer texture.
Additional results.
Fig. B shows the comparison with our ARC-NeRF, FreeNeRF [50], and FreeNeRF without the white and black prior. While FreeNeRF achieves high-quality rendering, it relies on a white and black prior, i.e. a highly heuristic approach considering the specific dataset’s characteristics. In contrast, our ARC-NeRF outperforms FreeNeRF without relying on such priors, even showing competitive performance compared to FreeNeRF with the prior.
The quantitative comparisons including the 6/9-view scenarios on DTU and more qualitative results are demonstrated in Tab. D and Fig. C, respectively. Note that we report the results of other methods from their original papers or [26, 35, 34], which outperformed the results from the corresponding original paper by modified training curriculum [26]. Our ARC-NeRF achieves competitive performance among the SOTA methods. Furthermore, our supplementary videos show the comparison with other methods on Realistic Synthetic 360 and FreeNeRF on DTU.
D Limitations and Future Work
Our proposed ARC-NeRF excels at capturing finer textures and details of object surfaces but shows limitations when addressing complex backgrounds, such as unbounded scenes composed of widely varying depths. This is due to potential obstacles between cast rays, which hinder pixel photo-consistency. Developing a new ray parameterization that can cover broad unseen view areas across significantly varied depths while effectively dealing with the obstacles would be a meaningful extension toward few-shot view synthesis for unbounded scenes as a future work.